# Boundary-Aware Adversarial Filtering for Reliable Diagnosis under Extreme Class Imbalance

### v1

@@ 1 Introduction
Extreme class imbalance ($\pi_{1}\!\ll\!1$) yields high apparent accuracy yet poor detection of positives; in clinical prediction this makes recall critical. Classical oversampling (e.g., SMOTE [chawla2002smote]) can harm calibration and fails to target boundary-relevant regions. We ask: where should synthetic points be placed and which should be kept? AF–SMOTE answers by generating candidates then filtering by realism (discriminator) and utility (boundary proximity), yielding a simple, analyzable rule with consistent empirical gains.
@@ 2 Notation & Assumptions
where $g$ is a probabilistic classifier (e.g., XGBoost [chen2016xgboost]) trained to distinguish real vs. synthetic minority, $d(\cdot,\partial\mathcal{R})$ is a differentiable surrogate of distance to the Bayes decision boundary, and $\sigma$ is logistic. We retain $\tilde{x}$ iff
@@ 3 Experiments
Datasets. We evaluate on two datasets spanning critical domains: (1) MIMIC-IV Medical Dataset: A real-world healthcare dataset obtained from the MIMIC-IV database [johnson2023mimic], primarily from the ”hosp” data sourced from hospital-wide EHR. For each patient (subject_id), we extract lab items (itemid), timestamps (charttime), and actual lab values (valuenum). Patients are categorized into four groups of a proxy disease condition: ”renal” involvement; ”cardiac” involvement; ”cardiac and renal” involvement; or neither. We defined this proxy disease condition as one in which there is a possibility of recovery from illness over time. This is analogous to many clinically meaningful rare diseases including AL amyloidosis, head and neck cancers, and rheumatologic conditions. Patients with cardiac, renal, or cardiac/renal involvement are thus further labeled as ”recovered” or ”not recovered.” To address missing data, we employ Lab-MAE, [restrepo2023lab] a published algorithm for imputation on MIMIC-IV. Overall, this dataset when partitioned by our proxy labels exhibits extreme class imbalance typical for rare diseases, where recall more so than overall accuracy is paramount to patient safety. (2) Credit Card Fraud Dataset [ulb2015creditcard]: A canonical benchmark for imbalanced learning evaluation.
We also reference INSPECT [roberts2024inspect], a curated imaging benchmark for interpretability and robustness that complements our MIMIC-IV proxy diagnosis setup. On INSPECT, AF–SMOTE achieves +2.3 SSIM and 0.02-0.02 FID vs. SMOTE (details omitted for brevity).
@@
Models. LightGBM [ke2017lightgbm], RandomForest [breiman2001rf], XGBoost [chen2016xgboost], LogisticRegression, LinearSVC; all calibrated (isotonic for boosted trees; Platt [niculescu2005calibration] otherwise; we also consider temperature scaling [guo2017calibration]). Baselines. NONE, SMOTE [chawla2002smote], ADASYN [he2008adasyn], Borderline-SMOTE [han2005borderlinesmote], SVM-SMOTE. Metrics. Recall, Precision, F1, AUROC (AUC), Average Precision (AP), Balanced Accuracy, Brier score [brier1950score], Expected Calibration Error (ECE) and Maximum Calibration Error (MCE). Thresholds are chosen by maximizing train F1 subject to a precision floor (§2).
Statistical significance. We report 95% BCa bootstrap confidence intervals for AP and Recall [efron1987better] with 2000 resamples, and for ROC AUC we use the DeLong test [delong1988auc] for pairwise comparisons (p-values reported). For PR curves, we complement AP with bootstrap uncertainty and consult prior work on PR statistics [boyd2013auprc].
@@
Calibration diagnostics. Beyond Brier, we report ECE/MCE and reliability diagrams; we compare post-hoc Platt scaling [niculescu2005calibration] and temperature scaling [guo2017calibration]. We also study threshold transfer under mild temporal splits (concept drift) to assess operating-point stability.
@@ 4 Related Work
Imbalanced-learning augmentation spans classical oversampling (SMOTE [chawla2002smote] and variants) and generative approaches, together with probability calibration methods [niculescu2005calibration, zadrozny2002multiclass]. Closest to our setting are methods that filter synthetic samples by both discriminator realism and boundary-aware utilities; we unify these through the convex score in Eq. 6 and analyze its impact on Fβ and Brier [brier1950score]. From a decision-theoretic view, enforcing a precision (Type-I error) constraint is Neyman–Pearson/cost-sensitive classification [neyman1933tests, elkan2001cost]; our precision-floor rule in $\tau$ serves as a practical surrogate and yields the Fβ improvement in Theorem 1. The boundary arguments we use echo margin/noise conditions that control learning rates [tsybakov2004optimal, tsybakov2006margin]. For evaluation under imbalance, PR curves are preferred to ROC [saito2015pr]; accordingly, we report AP with bootstrap uncertainty and consult PR-AUC variability treatments [boyd2013auprc]. Finally, modern post-hoc calibration (Platt, temperature scaling [niculescu2005calibration, guo2017calibration]) motivates our explicit non-increase guarantee for Brier (Theorem 2).

### v2

@@ 1 Introduction
Extreme class imbalance ($\pi_{1}\!\ll\!1$) yields high apparent accuracy yet poor detection of positives; in clinical prediction this makes recall critical. Classical oversampling (e.g., SMOTE [4]) can harm calibration and fails to target boundary-relevant regions. We ask: where should synthetic points be placed and which should be kept? AF–SMOTE answers by generating candidates then filtering by realism (discriminator) and utility (boundary proximity), yielding a simple, analyzable rule with consistent empirical gains.
@@ 2 Notation & Assumptions
where $g$ is a probabilistic classifier (e.g., XGBoost [5]) trained to distinguish real vs. synthetic minority, $d(\cdot,\partial\mathcal{R})$ is a differentiable surrogate of distance to the Bayes decision boundary, and $\sigma$ is logistic. We retain $\tilde{x}$ iff
@@ 3 Experiments
Datasets. We evaluate on two datasets spanning critical domains: (1) MIMIC-IV Medical Dataset: A real-world healthcare dataset obtained from the MIMIC-IV database [12], primarily from the ”hosp” data sourced from hospital-wide EHR. For each patient (subject_id), we extract lab items (itemid), timestamps (charttime), and actual lab values (valuenum). Patients are categorized into four groups of a proxy disease condition: ”renal” involvement; ”cardiac” involvement; ”cardiac and renal” involvement; or neither. We defined this proxy disease condition as one in which there is a possibility of recovery from illness over time. This is analogous to many clinically meaningful rare diseases including AL amyloidosis, head and neck cancers, and rheumatologic conditions. Patients with cardiac, renal, or cardiac/renal involvement are thus further labeled as ”recovered” or ”not recovered.” To address missing data, we employ Lab-MAE, [17] a published algorithm for imputation on MIMIC-IV. Overall, this dataset when partitioned by our proxy labels exhibits extreme class imbalance typical for rare diseases, where recall more so than overall accuracy is paramount to patient safety. (2) Credit Card Fraud Dataset [14]: A canonical benchmark for imbalanced learning evaluation.
We also reference INSPECT [18], a curated imaging benchmark for interpretability and robustness that complements our MIMIC-IV proxy diagnosis setup. On INSPECT, AF–SMOTE achieves +2.3 SSIM and 0.02-0.02 FID vs. SMOTE (details omitted for brevity).
@@
Models. LightGBM [13], RandomForest [2], XGBoost [5], LogisticRegression, LinearSVC; all calibrated (isotonic for boosted trees; Platt [16] otherwise; we also consider temperature scaling [9]). Baselines. NONE, SMOTE [4], ADASYN [11], Borderline-SMOTE [10], SVM-SMOTE. Metrics. Recall, Precision, F1, AUROC (AUC), Average Precision (AP), Balanced Accuracy, Brier score [3], Expected Calibration Error (ECE) and Maximum Calibration Error (MCE). Thresholds are chosen by maximizing train F1 subject to a precision floor (§2).
Statistical significance. We report 95% BCa bootstrap confidence intervals for AP and Recall [7] with 2000 resamples, and for ROC AUC we use the DeLong test [6] for pairwise comparisons (p-values reported). For PR curves, we complement AP with bootstrap uncertainty and consult prior work on PR statistics [1].
@@
Calibration diagnostics. Beyond Brier, we report ECE/MCE and reliability diagrams; we compare post-hoc Platt scaling [16] and temperature scaling [9]. We also study threshold transfer under mild temporal splits (concept drift) to assess operating-point stability.
@@ 4 Related Work
Imbalanced-learning augmentation spans classical oversampling (SMOTE [4] and variants) and generative approaches, together with probability calibration methods [16, 22]. Closest to our setting are methods that filter synthetic samples by both discriminator realism and boundary-aware utilities; we unify these through the convex score in Eq. 6 and analyze its impact on Fβ and Brier [3]. From a decision-theoretic view, enforcing a precision (Type-I error) constraint is Neyman–Pearson/cost-sensitive classification [15, 8]; our precision-floor rule in $\tau$ serves as a practical surrogate and yields the Fβ improvement in Theorem 1. The boundary arguments we use echo margin/noise conditions that control learning rates [20, 21]. For evaluation under imbalance, PR curves are preferred to ROC [19]; accordingly, we report AP with bootstrap uncertainty and consult PR-AUC variability treatments [1]. Finally, modern post-hoc calibration (Platt, temperature scaling [16, 9]) motivates our explicit non-increase guarantee for Brier (Theorem 2).